OpenStudy (anonymous):

how to complete the square? x^2+4x-7=0 explain how??

OpenStudy (solomonzelman):

\[ x^2+4x-7=0\]\[ x^2+4x=7\]\[ x^2+4x+4=7+4\]\[x^2+4x+4=11\]Go from here.

OpenStudy (anonymous):

so you take the middle, b, half it, then square it, then add it to both sides?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

 I find it is usually easiest to start on the second row. That way you are just dividing the deal in half. Before trying to add the deal (squared) back into your original equation.